The Florida Mining Company conducts recapitalization (or overhaul) operations on large, heavy mining equipment. Vehicles scheduled for periodic overhaul and repair are brought to the depot and a complete overhaul returns them back to the minim company in a “like new” condition. In order for operations to run smoothly, spare parts must be available so that the production line does not idle. Below is a table of three critical parts used in the overhaul process. The annual holding and backorder costs are given below. The backorder cost for each item is a one-time backorder cost (the cost does not accrue over time) and represents the cost of expediting that specific part. The holding cost is the cost of holding one unit of that particular part number in inventory for one year. Assume 52 weeks per year.  

Part Unit Cost Avg Weekly Demand Std Dev Weekly Demand Lead Time (in weeks) Current On-hand Inventory Ordering Cost Annual Holding Cost Backorder Cost
1001 $8.50 19 6 2 38 $125 $55 $1,000
1002 $47.25 7 5 16 112 $225 $45 $1,000
1003 $202.50 .50 .80 6 3 $175 $125 $1,000


Using the data from the Florida Mining Company, what is the target service level for part 1002 given the holding and backorder costs reported in the table for this item?

Answer Options:

.96

.98

.99

.85
Answer #1

Cost of holding = $45, This is cost of overstocking (Co) the part 1002.

Cost of backorder = $1000. This is the cost of understocking (Cu) the part 1002.

Critical ratio Service level = Cu / (Cu + Co) = 1000 / 1045 =0.9569 = 0.96